Q.19 The average cost of producing 10 units is Rs 30, while the average cost of producing 20 units is Rs 20. Find the average cost of producing 30 units.

{Hint : Find Fixed cost. TC of 10 units = Rs 300, TC of 20 units = Rs 400, VC = Rs 10/ unit

Fixed cost = TC = FC + VC (10* 10) = 300 = FC + 100 = FC = 200. Thus, AC of producing 30 units will be Rs 16.67}



1
Expert's answer
2021-05-30T14:19:19-0400

Given :

Average cost of producing 10 units = Rs 30

Average cost of producing 20 units = Rs 20.


Average cost "= \frac{Total\space cost } {Quantity}"

Total cost "= Average \space cost × Quantity"

Total cost at 10 units"= 30 \u00d7 10 = 300"

Total cost at 20 units "= 20 \u00d7 20 = 400"

Variable cost is that cost which changes with output.

Therefore variable cost = 10 per unit

"TC = TFC + TVC"

"TC = TFC + VC (Q)"

TC (at 10 units) = "TFC + 10(10) = 300"

"TFC = 200"

TC of producing 30 units "= TFC + TVC = 200 + 10(30) = 200 + 300 = 500"

AC of producing 30 units"= \frac{500}{30}"

"= 16.66"